the pineapple juice comes in two concentrated forms, 17% and 30% pineapple juice, which must be combines into a solution that can be mixed with the orange juice. if 50 gallons of 17% juice is used, how many gallons of the 30% juice must be used to obtain 20% pineapple solution
Answer by mananth(16949)

--------percent ---------------- quantity
Juice I 17 ---------------- 15 gallons
juice II 30 ---------------- x gallons
Mixture 20 ---------------- 15 + x gallons

17 * 15 + 30 x = 20 ( 15 + x )
255 + 30 x = 300 + 20 x
30 x + -20 x = 300 - -255
10 x = 45
/ 10
x = 4.5
4.5 gallons of 30 % juice II
m.ananth@hotmail.ca